OpenStudy (anonymous):

what is the solution to the proportion 12/25=8/x

OpenStudy (anonymous):

x=16 and 2/3

OpenStudy (texaschic101):

do you know how to cross multiply ?

OpenStudy (texaschic101):

12/25 = 8/x (cross multiply) (12)(x) = (25)(8) 12x = 200 x = 200/12 x = 16 8/12 which reduces to 16 2/3

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

yw